You found a treasure map,and it's said: Staring at well walk 30m north, 20m 30°N of East, 15m 18° south of west,and finally 10m north. How far is the treasure map from the well?


1
Expert's answer
2021-09-30T18:28:16-0400

The distance is


"R_N=30+20\sin30°-15\sin18°+10=45.4\text{ m},\\
R_E=20\cos30°-15\cos18°=-4.27\text{ m},\\

R=\sqrt{(R_N)^2+(R_E)^2}=46\text{ m}."